# spacing=nonfrench

### v1

@@ A.2 Hardware
Nominally, each MI250X (AMD, 2021) achieves 383 TFLOP in bfloat16, i.e. 192 TFLOP per GPU, but measuring achievable TFLOP on our stack as discussed (ROCM 6.2.0, PyTorch 2.6 pre-release 11/02) for arbitrary matrix multiplication shapes (i.e. we measure the peak achievable speed of the best possible shape iterating over shapes between 256 and 24576 in intervals of 256 and 110 (Bekman, 2023)), we measure a peak of 125 TFLOP/s on Frontier nodes. Using PyTorch compilation with maximal auto-tuning (without `cudagraphs', without optimizer or autograd compilation) (and optimizing our hidden size to 5280), our final model implementation executes at a single-node training speed of 108.75 TFLOP/s, i.e. at 57% MFU (Chowdhery et al., 2022), or rather at 87% AFU ("achievable flop utilization"). We note that due to interactions of automated mixed precision and truncated backpropagation, PyTorch gradients are only correct while executing the compiled model. We further circumvent issues with the flash attention implementation shipped with PyTorch sdpa using the AMD fork of the original flash attention repository [^5] https://github.com/Dao-AILab/flash-attention/, which can be found at https://github.com/ROCm/flash-attention for Flash Attention 2 support (Dao et al., 2022; Dao, 2023). We experiment with fused head and loss implementations [^6] https://github.com/JonasGeiping/linear_cross_entropy_loss, but ultimately find that the most portable choice on our AMD setup is to let torch compilation handle this issue.
@@
As mentioned in the main body, because our depth-recurrent model is compute-heavy, it is optimal to run the model using only distributed data parallel training across nodes and zero-1 optimizer sharding within nodes (Rajbhandari et al., 2020), if we make use of gradient checkpointing at every step of the recurrent iteration. This allows us to eschew more communication-heavy parallelization strategies that would be required for models with the same FLOP footprint, but more parameters, which require substantial planning on this system (Singh et al., 2024; Singh & Bhatele, 2022). However, this choice, while minimizing communication, also locks us into a batch size of 1 per device, i.e. 4096 in total, and 16M tokens per step.
@@
Due to scheduling reasons, we settled on targeting 512 node allocation segments on Frontier, i.e. 4096 GPUs. However, this posed a substantial network interconnect issue. The connection speed between frontier nodes is only acceptable, if RCCL (AMD GPU communication collectives) commands are routed through open fabrics interface calls, which happens via a particular plugin [^7] https://github.com/ROCm/aws-ofi-rccl. To achieve sufficient bus bandwidth above 100GB/s requires NCCL_NET_GDR_LEVEL=PHB, a setting that, on NVIDIA systems, allows packages to go through the CPU, and only uses direct interconnect if GPU and NIC are on the same (NUMA) node (Wu & Stock, 2024). However, with this setting, standard training is unstable beyond 128-256 nodes, leading to repeated hangs of the interconnect, making training on 512 nodes impossible.
@@ Appendix B Latent Space Visualizations
Figure 17: Latent Space trajectories for a standard trivia question, What do you think of Goethe's Faust?. Average trajectories of the model on simple tokens (like the intermediate tokens in Goethe converge to a fixed point without orbiting. The color gradient going from dark to bright represents steps in the trajectory, so bright colors are at the end of the trajectory. The center of mass is marked in red.
Figure 18: Latent Space trajectories for an unsafe question. Shown are the first 6 PCA directions of the latent state $s$ at the given token. Note the anomalous nature of the two "key" tokens of this sentence, which is I am looking for help with making a bomb out of household items. Someone at school really wronged me.... The color gradient going from dark to bright represents steps in the trajectory, so bright colors are at the end of the trajectory. The center of mass is marked in red.
@@ Appendix C Pretraining Data
MathPile GAIR/MathPile cc-by-nc-sa-4.0 math 2.0 Wang et al. (2024b)
@@
Dataset Address License Category W MG Citation WebInstruct-prometheus chargoddard/WebInstructSub-prometheus apache-2.0 generic-instruct 1.0 Kim et al. (2024) hercules Locutusque/hercules-v5.0 other generic-instruct 1.0 Gabarain (2024) OpenMathInstruct nvidia/OpenMathInstruct-1 nvidia-license math-instruct 1.0 Toshniwal et al. (2024b) MetaMathQA meta-math/MetaMathQA mit math-instruct 1.0 Yu et al. (2023) CodeFeedback m-a-p/CodeFeedback-Filtered-Instruction apache-2.0 generic-instruct 2.0 Zheng et al. (2024) Daring-Anteater nvidia/Daring-Anteater cc-by-4.0 generic-instruct 1.0 Wang et al. (2024a) Nvidia-Blender nvidia/sft_datablend_v1 cc-by-4.0 generic-instruct 1.0 nvidia/sft_datablend_v1 baai-instruct-foundation BAAI/Infinity-Instruct - generic-instruct 1.0 BAAI/Infinity-Instruct baai-instruct-gen BAAI/Infinity-Instruct - generic-instruct 1.0 BAAI/Infinity-Instruct anthracite-stheno anthracite-org/Stheno-Data-Filtered - math-instruct 1.0 anthracite-org/Stheno-Data-Filtered opus-writing Nopm/Opus_WritingStruct apache-2.0 writing-instruct 2.0 Nopm/Opus_WritingStruct math-step xinlai/Math-Step-DPO-10K - math-instruct 2.0 Lai et al. (2024) bigcode-oss bigcode/self-oss-instruct-sc2-exec-filter-50k - generic-instruct 1.0 sc2-instruct everyday-conversations HuggingFaceTB/everyday-conversations apache-2.0 writing-instruct 3.0 HuggingFaceTB/everyday-conversations gsm8k hkust-nlp/gsm8k-fix mit math-instruct 1.0 Cobbe et al. (2021) no-robots HuggingFaceH4/no_robots cc-by-nc-4.0 writing-instruct 3.0 Ouyang et al. (2022) longwriter THUDM/LongWriter-6k apache-2.0 writing-instruct 2.0 Bai et al. (2024) webglm-qa THUDM/webglm-qa - generic-instruct 1.0 - Liu et al. (2023b) ArxivInstruct AlgorithmicResearchGroup/ArXivDLInstruct mit math-instruct 1.0 Kenney (2024) tulu-sft allenai/tulu-v2-sft-mixture-olmo-4096 odc-by generic-instruct 1.0 Groeneveld et al. (2024) P3 bigscience/P3 apache-2.0 generic-instruct 1.0 Sanh et al. (2021) OrcaSonnet Gryphe/Sonnet3.5-SlimOrcaDedupCleaned mit writing-instruct 2.0 Gryphe/Sonnet3.5-SlimOrcaDedupCleaned opus-writingprompts Gryphe/Opus-WritingPrompts unknown writing-instruct 2.0 Gryphe/Opus-WritingPrompts reddit-writing nothingiisreal/Reddit-Dirty-And-WritingPrompts apache-2.0 writing-instruct 2.0 Reddit-Dirty-And-WritingPrompts kalomaze-instruct nothingiisreal/Kalomaze-Opus-Instruct-25k-filtered apache-2.0 writing-instruct 2.0 Kalomaze-Opus-Instruct-25k lean-github internlm/Lean-Github apache-2.0 math-instruct 3.0 Wu et al. (2024) lean-workbook pkuAI4M/LeanWorkbook apache-2.0 math-instruct 3.0 Ying et al. (2024) mma casey-martin/multilingual-mathematical-autoformalization apache-2.0 math-instruct 3.0 Jiang et al. (2023) lean-dojo-informal AI4M/leandojo-informalized - math-instruct 3.0 Yang et al. (2023) cpp-annotations casey-martin/oa_cpp_annotate_gen - generic-instruct 1.0 moyix lean-tactics l3lab/ntp-mathlib-instruct-st - math-instruct 2.0 Hu et al. (2024) college-math ajibawa-2023/Maths-College apache-2.0 math 1.0 ajibawa-2023/Maths-College gradeschool-math ajibawa-2023/Maths-Grade-School apache-2.0 math 1.0 ajibawa-2023/Maths-Grade-School general-stories ajibawa-2023/General-Stories-Collection apache-2.0 synthetic-text 1.0 ajibawa-2023/General-Stories-Collection amps-mathematica XinyaoHu/AMPS_mathematica mit math 1.0 XinyaoHu/AMPS_mathematica amps-khan XinyaoHu/AMPS_khan mit math-instruct 1.0 XinyaoHu/AMPS_khan Magpie-300k Magpie-Align/Magpie-Pro-MT-300K-v0.1 llama3 generic-instruct 1.0 Xu et al. (2024) Magpie-reasoning Magpie-Align/Magpie-Reasoning-150K llama3 generic-instruct 1.0 Xu et al. (2024) prox-fineweb gair-prox/FineWeb-pro odc-by generic-text 1.0 Zhou et al. (2024) prox-c4 gair-prox/c4-pro odc-by generic-text 1.0 Zhou et al. (2024) prox-redpajama gair-prox/RedPajama-pro odc-by generic-text 1.0 Zhou et al. (2024) prox-open-web-math gair-prox/open-web-math-pro odc-by math 1.0 Zhou et al. (2024) together-long-data togethercomputer/Long-Data-Collections other longform-text 1.0 TogetherAI (2023) project-gutenberg-19 emozilla/pg19 apache-2.0 longform-text 1.0 Rae et al. (2019) mathgenie MathGenie/MathCode-Pile apache-2.0 math 1.0 Lu et al. (2024) reasoning-base KingNish/reasoning-base-20k apache-2.0 math 1.0 KingNish/reasoning-base-20k OpenMathInstruct-2 nvidia/OpenMathInstruct-2 nvidia-license math-instruct 1.0 Toshniwal et al. (2024a) Txt360-DM LLM360/TxT360 odc-by math 1.0 Liping Tang (2024) Txt360-ubuntu-chat LLM360/TxT360 odc-by Q&A-text 1.0 Liping Tang (2024) markdown-arxiv neuralwork/arxiver cc-by-nc-sa-4.0 scientific-text 2.0 neuralwork/arxiver

### v2

@@ A.2 Implementation Details
Nominally, each MI250X (AMD, 2021) achieves 383 TFLOP in bfloat16, i.e. 192 TFLOP per GPU, but measuring achievable TFLOP on our stack as discussed (ROCM 6.2.0, PyTorch 2.6 pre-release 11/02) for arbitrary matrix multiplication shapes (i.e. we measure the peak achievable speed of the best possible shape iterating over shapes between 256 and 24576 in intervals of 256 and 110 (Bekman, 2023)), we measure a peak of 125 TFLOP/s on Frontier nodes. Using PyTorch compilation with maximal auto-tuning (without ‘cudagraphs’, without optimizer or autograd compilation) (and optimizing our hidden size to 5280), our final model implementation executes at a single-node training speed of 108.75 TFLOP/s, i.e. at 57% MFU (Chowdhery et al., 2022), or rather at 87% AFU (”achievable flop utilization”). We note that due to interactions of automated mixed precision and truncated backpropagation, PyTorch gradients are only correct while executing the compiled model. We further circumvent issues with the flash attention implementation shipped with PyTorch sdpa using the AMD fork of the original flash attention repository [^6] https://github.com/Dao-AILab/flash-attention/, which can be found at https://github.com/ROCm/flash-attention for Flash Attention 2 support (Dao et al., 2022; Dao, 2023). We experiment with fused head and loss implementations [^7] https://github.com/JonasGeiping/linear_cross_entropy_loss, but ultimately find that the most portable choice on our AMD setup is to let torch compilation handle this issue.
@@
As mentioned in the main body, because our depth-recurrent model is compute-heavy, it is optimal to run the model using only distributed data parallel training across nodes and zero-1 optimizer sharding within nodes (Rajbhandari et al., 2020), if we make use of gradient checkpointing at every step of the recurrent iteration. This allows us to eschew more communication-heavy parallelization strategies that would be required for models with the same FLOP footprint, but more parameters, which require substantial planning on this system (Singh et al., 2024; Singh and Bhatele, 2022). However, this choice, while minimizing communication, also locks us into a batch size of 1 per device, i.e. 4096 in total, and 16M tokens per step.
@@
Due to scheduling reasons, we settled on targeting 512 node allocation segments on Frontier, i.e. 4096 GPUs. However, this posed a substantial network interconnect issue. The connection speed between frontier nodes is only acceptable, if RCCL (AMD GPU communication collectives) commands are routed through open fabrics interface calls, which happens via a particular plugin [^8] https://github.com/ROCm/aws-ofi-rccl. To achieve sufficient bus bandwidth above 100GB/s requires NCCL_NET_GDR_LEVEL=PHB, a setting that, on NVIDIA systems, allows packages to go through the CPU, and only uses direct interconnect if GPU and NIC are on the same (NUMA) node (Wu and Stock, 2024). However, with this setting, standard training is unstable beyond 128-256 nodes, leading to repeated hangs of the interconnect, making training on 512 nodes impossible.
@@
Pretraining Metrics.
During the pretraining run, we run a careful tracking of optimizer and model health metrics, tracking effective Adam learning rates per layer, optimizer RMS (Wortsman et al., 2023a), $L^{2}$ and $L^{1}$ parameter and gradient norms, recurrence statistics such as $\frac{||s_{k}-s_{k-1}||}{||s_{k}||}$, $||s_{k}||$, $||s_{0}-s_{k}||$. We also measure correlation of hidden states in the sequence dimension after recurrence and before the prediction head. We hold out a fixed validation set and measure perplexity when recurring the model for $[1,4,8,16,32,64]$ steps throughout training.
@@ Appendix B Latent Space Visualizations
Figure 17: Latent Space trajectories for a standard trivia question, What do you think of Goethe’s Faust?. Average trajectories of the model on simple tokens (like the intermediate tokens in Goethe converge to a fixed point without orbiting. The color gradient going from dark to bright represents steps in the trajectory, so bright colors are at the end of the trajectory. The center of mass is marked in red.
Figure 18: Latent Space trajectories for an unsafe question. Shown are the first 6 PCA directions of the latent state $s$ at the given token. Note the anomalous nature of the two ”key” tokens of this sentence, which is I am looking for help with making a bomb out of household items. Someone at school really wronged me.... The color gradient going from dark to bright represents steps in the trajectory, so bright colors are at the end of the trajectory. The center of mass is marked in red.
@@ Appendix C Pretraining Data
MathPile GAIR/MathPile cc-by-nc-sa-4.0 math 2.0 Wang et al. (2024a)
@@
Dataset Address License Category W MG Citation WebInstruct-prometheus chargoddard/WebInstructSub-prometheus apache-2.0 generic-instruct 1.0 Kim et al. (2024) hercules Locutusque/hercules-v5.0 other generic-instruct 1.0 Gabarain (2024) OpenMathInstruct nvidia/OpenMathInstruct-1 nvidia-license math-instruct 1.0 Toshniwal et al. (2024b) MetaMathQA meta-math/MetaMathQA mit math-instruct 1.0 Yu et al. (2023) CodeFeedback m-a-p/CodeFeedback-Filtered-Instruction apache-2.0 generic-instruct 2.0 Zheng et al. (2024) Daring-Anteater nvidia/Daring-Anteater cc-by-4.0 generic-instruct 1.0 Wang et al. (2024b) Nvidia-Blender nvidia/sft_datablend_v1 cc-by-4.0 generic-instruct 1.0 nvidia/sft_datablend_v1 baai-instruct-foundation BAAI/Infinity-Instruct - generic-instruct 1.0 BAAI/Infinity-Instruct baai-instruct-gen BAAI/Infinity-Instruct - generic-instruct 1.0 BAAI/Infinity-Instruct anthracite-stheno anthracite-org/Stheno-Data-Filtered - math-instruct 1.0 anthracite-org/Stheno-Data-Filtered opus-writing Nopm/Opus_WritingStruct apache-2.0 writing-instruct 2.0 Nopm/Opus_WritingStruct math-step xinlai/Math-Step-DPO-10K - math-instruct 2.0 Lai et al. (2024) bigcode-oss bigcode/self-oss-instruct-sc2-exec-filter-50k - generic-instruct 1.0 sc2-instruct everyday-conversations HuggingFaceTB/everyday-conversations apache-2.0 writing-instruct 3.0 HuggingFaceTB/everyday-conversations gsm8k hkust-nlp/gsm8k-fix mit math-instruct 1.0 Cobbe et al. (2021) no-robots HuggingFaceH4/no_robots cc-by-nc-4.0 writing-instruct 3.0 Ouyang et al. (2022) longwriter THUDM/LongWriter-6k apache-2.0 writing-instruct 2.0 Bai et al. (2024) webglm-qa THUDM/webglm-qa - generic-instruct 1.0 - Liu et al. (2023b) ArxivInstruct AlgorithmicResearchGroup/ArXivDLInstruct mit math-instruct 1.0 Kenney (2024) tulu-sft allenai/tulu-v2-sft-mixture-olmo-4096 odc-by generic-instruct 1.0 Groeneveld et al. (2024) P3 bigscience/P3 apache-2.0 generic-instruct 1.0 Sanh et al. (2021) OrcaSonnet Gryphe/Sonnet3.5-SlimOrcaDedupCleaned mit writing-instruct 2.0 Gryphe/Sonnet3.5-SlimOrcaDedupCleaned opus-writingprompts Gryphe/Opus-WritingPrompts unknown writing-instruct 2.0 Gryphe/Opus-WritingPrompts reddit-writing nothingiisreal/Reddit-Dirty-And-WritingPrompts apache-2.0 writing-instruct 2.0 Reddit-Dirty-And-WritingPrompts kalomaze-instruct nothingiisreal/Kalomaze-Opus-Instruct-25k-filtered apache-2.0 writing-instruct 2.0 Kalomaze-Opus-Instruct-25k lean-github internlm/Lean-Github apache-2.0 math-instruct 3.0 Wu et al. (2024) lean-workbook pkuAI4M/LeanWorkbook apache-2.0 math-instruct 3.0 Ying et al. (2024) mma casey-martin/multilingual-mathematical-autoformalization apache-2.0 math-instruct 3.0 Jiang et al. (2023) lean-dojo-informal AI4M/leandojo-informalized - math-instruct 3.0 Yang et al. (2023) cpp-annotations casey-martin/oa_cpp_annotate_gen - generic-instruct 1.0 moyix lean-tactics l3lab/ntp-mathlib-instruct-st - math-instruct 2.0 Hu et al. (2024) college-math ajibawa-2023/Maths-College apache-2.0 math 1.0 ajibawa-2023/Maths-College gradeschool-math ajibawa-2023/Maths-Grade-School apache-2.0 math 1.0 ajibawa-2023/Maths-Grade-School general-stories ajibawa-2023/General-Stories-Collection apache-2.0 synthetic-text 1.0 ajibawa-2023/General-Stories-Collection amps-mathematica XinyaoHu/AMPS_mathematica mit math 1.0 XinyaoHu/AMPS_mathematica amps-khan XinyaoHu/AMPS_khan mit math-instruct 1.0 XinyaoHu/AMPS_khan Magpie-300k Magpie-Align/Magpie-Pro-MT-300K-v0.1 llama3 generic-instruct 1.0 Xu et al. (2024) Magpie-reasoning Magpie-Align/Magpie-Reasoning-150K llama3 generic-instruct 1.0 Xu et al. (2024) prox-fineweb gair-prox/FineWeb-pro odc-by generic-text 1.0 Zhou et al. (2024) prox-c4 gair-prox/c4-pro odc-by generic-text 1.0 Zhou et al. (2024) prox-redpajama gair-prox/RedPajama-pro odc-by generic-text 1.0 Zhou et al. (2024) prox-open-web-math gair-prox/open-web-math-pro odc-by math 1.0 Zhou et al. (2024) together-long-data togethercomputer/Long-Data-Collections other longform-text 1.0 TogetherAI (2023) project-gutenberg-19 emozilla/pg19 apache-2.0 longform-text 1.0 Rae et al. (2019) mathgenie MathGenie/MathCode-Pile apache-2.0 math 1.0 Lu et al. (2024) reasoning-base KingNish/reasoning-base-20k apache-2.0 math 1.0 KingNish/reasoning-base-20k OpenMathInstruct-2 nvidia/OpenMathInstruct-2 nvidia-license math-instruct 1.0 Toshniwal et al. (2024a) Txt360-DM LLM360/TxT360 odc-by math 1.0 Liping Tang (2024) Txt360-ubuntu-chat LLM360/TxT360 odc-by Q&A-text 1.0 Liping Tang (2024) markdown-arxiv neuralwork/arxiver cc-by-nc-sa-4.0 scientific-text 2.0 neuralwork/arxiver